OpenStudy (anonymous):

1. 2t + 8 >= -4(t + 1) 2. 4y + 3 <= y + 6 I know the answers ( t >= -2, y <= 1 ) I just need to know the steps.

OpenStudy (texaschic101):

2t + 8 >= -4(t + 1) -- distribute through the parenthesis 2t + 8 >= -4t - 4 -- add 4t to both sides 2t + 4t + 8 >= -4 -- subtract 8 from both sides 2t + 4t >= -4 - 8 -- combine like terms 6t >= -12 -- divide both sides by 6 t >= -2 4y + 3 <= y + 6 -- subtract y from both sides 4y - y + 3<= 6 -- subtract 3 from both sides 4y - y <= 6 - 3 -- combine like terms 3y <= 3 -- divide both sides by 3 y <= 1

OpenStudy (anonymous):

@texaschic101 Sometimes get confused when there's a variable on both sides :P Thanks so much! :)

OpenStudy (texaschic101):

It can be confusing...glad I could help